Red Hawk Enterprises sells handmade clocks. Its variable cost per clock is $4.80, and each clock sells for $16.00.

Calculate Red Hawk's unit contribution margin.

Calculate Red Hawk's contribution margin ratio.

Suppose Red Hawk sells 1,700 clocks this year. Calculate the total contribution margin.
Answer #1

Contribution margin=Sales-Variable cost  

=(16-4.8)=$11.2 per unit

Contribution margin ratio=Contribution margin/Sales

=(11.2/16)=70%(or 0.7)

Total contribution margin would be =($11.2/unit*1700 clocks)

=$19040